please prepare the cash flow statement
Rhodes Corporation: Income Statements for Year Ending December 31 (Millions of Dollars) 2009 S10,000 8,500 360 S 1,140 100 040 416 2010 Sales $11,000 Operating costs excluding depreciation Depreciation Earnings before interest and taxes 9,360 Less interest Earnings before taxes S 1,260 120 S 1,140 S 1, Taxes (40%) Net income available to common stockholders S 684 S 220 Common dividends S 200 Rhodes Corporation: Bolance Sheets as of December 31 (Millions of Dollors) 2010 2009 Assets Cash Short-term investments Accounts receivable Inventories Total current assets Net plant and equipment Total assets S 550 110 2,750 1,650 S5,060 S 500 100 2,500 1,500 $4,600 . 3.500 $8,910 Liabilities and Equity Accounts payable Accruals Notes payable Total current liabilities Long-term debt Total liabilities Common stock Retained earnings Total common equity Total liabilities and equity $1,100 550 S1,000 500 200 S1,700 1,000 $2,700 4,400 $2,034 S3,134 4,312 1,464 S5,400 $8,910
Answer #1
STATEMENT OF CASH FLOWS
Cash flows from operating activities:
Net income 684
Adjustments to reconcile net income to net cash profived by operations:
Depreciation expense 380
Increase in short term investments (110-100) -10
Increase in accounts receivable = 2750-2500 = -250
Increase in inventory = 1650-1500 = -150
Increase in income accruals = 550-500 = 50
Increase in accounts payable = 1100-1000 = 100 120
Net cash provided by operating activities 804
Cash flows from investing activities:
Purchase of plant equipment (3850-3500+380) -730
Net cash used in investing activities -730
Cash flows from financing activities:
Cash from notes payable (384-200) 184
Purchase of treasury common stock (4400-4312) -88
Cash from long term debt (1100-1000) 100
Dividend (684-1464+1000) -220
Net cash used in financing activities -24
Net increase (decrease) in cash 50
Cash balance at beginning of year 500
Cash balance at end of year 550